please may someone help me i cant get this question right? tracey is five times as old as sipho. the sum of their ages is eighteen. how old are tracey and sipho
Found 3 solutions by Theo, Alan3354, ikleyn:
Answer by Theo(13342)

t = tracey's age.
s= sipho's age.
t = 5 * s
t + s = 18
replace t with 5 * s to get 5 * s + s =18
combine like terms to get 6 * s = 18
solve for s to get s = 18 / 6 = 3
that makes t = 5 * 3 = 15
t = 15 years of age.
s = 3 years of age.
t = 5 * s becomes 18 = 5 * 3 which becomes 18 = 18 which is true.
t + s = 18 becomes 15 + 3 = 18 which becomes 18 = 18 which is true.
solution is confirmed to be good.
solution is tracey is 15 years old and sipho is 3 years old.

Answer by Alan3354(69443)

Getting the question right is easy.
Getting the answer right can be more difficult.
-----------------
It would be educational (for you) if you showed what you tried.

Answer by ikleyn(52794)

.

Let " x " be the Sipho's age.


Then the Tracey's age is 5x, according to the condition.


The sum of their ages is 18.  It means that


    x + 5x = 18,    or

    6x     = 18.


Hence,   x = 18/6 = 3 years.   It is the Sipho's age.


Then the Tracey age is 5 times it, i.e.  5*3 = 15 years.


ANSWER.  Sipho is 3 years old.   Tracey is 15 years old.


Is it clear to you now ?